OpenStudy (anonymous):

Write the equation in standard form of the line through the point (4,-5) with a slope of -2/5.

OpenStudy (zale101):

\[y-y_1=m(x-x1)\]

OpenStudy (zale101):

a point on the line (x1, y1), and the slope is m

OpenStudy (anonymous):

y=mx+b -5=(-2/5)(4)+b solve for b

OpenStudy (anonymous):

-5=-8/5 +b what do I do next?

OpenStudy (anonymous):

@quentinloy

OpenStudy (anonymous):

so you would add 8/5 to both sides to solve for b; then you would write the equation as y=(-2/5)x-(17/5)

OpenStudy (anonymous):

I thought that when y is by itself it is slope intercept form?

OpenStudy (anonymous):

@quentinloy

OpenStudy (anonymous):

oh, you're right. I'm sorry. so you would add (2/5)x to both sides. (2/5)x+y=(-17/5)

OpenStudy (anonymous):

\[(2/5)x+y=(-17/5)\]